OpenStudy (anonymous):

condense to one logarithm: 1/2logz+5logx. Would it be sqrtlogz/logx^5?

Directrix (directrix):

(1/2)* logz + 5logx = log((z)^(1/2)) + log (x)^5 = log[ x^5 * sqr(z) ] I did not get what you did @yoitsren

Directrix (directrix):